What is the pH of the resultant solution if the pKa of an amine is 7.4 and the ratio of nonionized to ionized species is 10:1?
Answer #1

The pH can be determined using Henderson-Heasselbalch equation :

pH = pKa + log[A-]/[HA]

= 7.4 + log(1/10)

= 7.4 -1

= 6.4